# On finite group scheme-theoretical categories, II

### v1

@@ Abstract.
Let ${\mathscr{C}}:={\mathscr{C}}(G,\omega,H,\psi)$ be a finite group scheme-theoretical category over an algebraically closed field of characteristic $p\geq 0$ [G1]. For any indecomposable exact module category over ${\mathscr{C}}$, we classify its simple objects and provide an expression for their projective covers, in terms of double cosets and projective representations of certain closed subgroup schemes of $G$. This upgrades a result of Ostrik [O] for group-theoretical fusion categories in characteristic 00, and generalizes our previous work [GS] for the case $\omega=1$. As a byproduct, we describe the simples and indecomposable projectives of ${\mathscr{C}}$. Finally, we apply our results to describe the blocks of the center of ${\rm Coh}(G,\omega)$.
@@ 1. Introduction
Recall that group-theoretical fusion categories [O] play a fundamental role in the theory of fusion categories in characteristic 00 (see [GS, Section 1] and references therein), which suggests that finite group scheme-theoretical categories in characteristic $p>0$ play an important role in the theory of finite tensor categories, thus making them interesting tensor categories worthwhile to study.
@@
Unlike in characteristic 00, finite group scheme-theoretical categories and their module categories are rarely semisimple if $p>0$, and a plethora of questions arises. In [GS], we addressed some of these questions in the case $\omega=1$, and the goal of this work is to extend that treatment to arbitrary $\omega\in Z^{3}(G,\mathbb{G}_{m})$.
@@
${\rm Fun}_{{\rm Coh}(G,\omega)}\left(\mathscr{M}(H,\psi),\mathscr{M}(K,\eta)\right)\cong{\mathscr{M}}\left((H,\psi),(K,\eta)\right),$
@@
Finally, define (see Definition 2.7) the abelian categories
$\mathscr{A}:=\mathrm{Coh}^{\left(\left(B,\psi\right),\left(C,\eta\right)\right)}(A\backslash(B\times C),\Phi),\,\,\,{\rm and}$
$\mathscr{B}:=\mathrm{Coh}^{\left(\left(A\times B,\xi^{-1}\times\psi\right),\left(C,\eta\right)\right)}(B\times C,\Phi).$
@@
[…]
@@
${\rm F}\colon{\rm Corep}_{k}(A,\xi^{-1})\xrightarrow{\cong}\mathrm{Coh}^{\left(\left(B,\psi\right),\left(C,\eta\right)\right)}(A\backslash(B\times C),\Phi),$
@@
    $\mathbf{Ind}_{Z}:{\rm Corep}_{k}(L^{g},\xi_{g}^{-1})\xrightarrow{\cong}{\mathscr{M}}_{Z},$
@@
    There is a bijection between equivalence classes of pairs $(Z,V)$, where $Z\in Y(k)$ is a closed point with representative $g\in Z(k)$, and $V\in{\rm Corep}_{k}(L^{g},\xi_{g}^{-1})$ is simple, and simple objects of $\mathscr{M}$, assigning $(Z,V)$ to $\mathbf{Ind}_{Z}(V)$.
@@
    $\mathbf{F}_{Z}:{\rm Corep}_{k}(L^{g},\xi_{g}^{-1})\xrightarrow{\cong}{\mathscr{M}}_{Z}.$
@@
$\mathbf{F}_{C}:{\rm Rep}_{k}(G_{C},\omega_{g})\xrightarrow{\cong}\mathscr{Z}(G,\omega)_{C},\,\,\,V\mapsto\mathscr{O}(C)\otimes_{k}V,$
@@
$F:\mathscr{Z}(G,\omega)\xrightarrow{\cong}\mathscr{Z}_{\mathscr{D}^{\circ}}\left(\mathscr{D}\right)^{G(k)}=\left(\bigoplus_{a\in G(k)}\mathscr{Z}_{\mathscr{D}^{\circ}}\left(\mathscr{D}^{\circ}\boxtimes a\right)\right)^{G(k)}.$
@@
$F_{C}:\overline{\mathscr{Z}(G,\omega)_{C}}\xrightarrow{\cong}\bigoplus_{a\in C(k)}\mathscr{Z}_{\mathscr{D}^{\circ}}\left(\mathscr{D}^{\circ}\boxtimes a\right)^{G(k)}.$
@@
$F_{1}:\overline{{\rm Rep}_{k}(G)}\xrightarrow{\cong}\mathscr{Z}\left(G^{\circ},\omega^{\circ}\right)^{G(k)}.$
@@ 2. Preliminaries
We work over an algebraically closed field $k$ of characteristic $p\geq 0$. All schemes considered in this paper are assumed to be finite over $k$, unless otherwise stated. We assume familiarity with the theory of finite tensor categories and their modules categories, and refer to [EGNO] for any unexplained notion.
@@ 2.1. Sheaves on finite schemes
For a finite scheme $X$, let ${\rm Coh}(X)$ be the abelian category of sheaves on $X$, i.e., the category of finite dimensional representations of the finite dimensional commutative algebra $\mathscr{O}(X)$. Given a closed point $x\in X(k)$, let ${\rm Coh}(X)_{x}\subset{\rm Coh}(X)$ be the abelian subcategory of sheaves on $X$ supported on $x$, so
@@
Let $Y$ be a finite scheme, $\varphi:Y\to X$ a scheme morphism, and $\varphi^{\sharp}:\mathscr{O}(X)\to\mathscr{O}(Y)$ the corresponding algebra homomorphism. Recall that $\varphi$ induces a pair $(\varphi^{*},\varphi_{*})$ of adjoint functors of abelian categories
@@ 2.2. Finite group schemes
A finite group scheme $G$ is a finite scheme whose coordinate algebra $\mathscr{O}(G)$ is a (finite dimensional, commutative) Hopf algebra (see, e.g., [J, W]). It is called étale if $G=G(k)$, where $G(k)$ is the group of closed points of $G$; these finite group schemes are in correspondence with finite abstract groups (see, e.g., [W, Section 6.4]). On the other hand, $G$ is connected if $G(k)=1$. For example, if $\mathfrak{g}$ is a finite dimensional $p$-Lie algebra over $k$ ($p>0$), its $p$-restricted universal enveloping algebra $u(\mathfrak{g})$ is a finite dimensional cocommutative Hopf algebra (see, e.g., [J]), and the dual commutative Hopf algebra $u(\mathfrak{g})^{*}$ is local, and satisfies $x^{p}=0$ for all $x$ in the augmentation ideal. Thus, the associated group scheme is connected. Moreover, any finite connected group scheme is obtained by successive extensions of dualized $p$-restricted enveloping algebras (see [DG, Section 2.7]).
@@
[…]
where $G^{\circ}$ is connected and $G(k)$ is étale. In particular, $G(k)$ acts on ${\rm Coh}(G)_{1}\cong{\rm Coh}(G^{\circ})$, say via $a\mapsto T_{a}$, and we have
(2.4) ${\rm Coh}(G)\cong{\rm Coh}(G^{\circ})\rtimes G(k).$
@@ 2.3. Quotients and free actions
Let $C$ be a finite group scheme, and $X$ a finite scheme equipped with a right $C$-action
@@
is a closed immersion. Recall [Mu, Theorem 1(B), p.112] that in this case, the morphism $(p_{1},\mu)$ induces a scheme isomorphism
@@
$f\otimes_{\mathscr{O}(X/C)}\tilde{f}\mapsto(f\otimes 1)\mu^{\sharp}(\tilde{f}).$
@@
Consider the obvious projections
$p_{1}:X\times C\twoheadrightarrow X,\,\,\,\text{p}_{1}:X\times C\times C\twoheadrightarrow X,$
${\rm and}\,\,\,p_{12}:X\times C\times C\twoheadrightarrow X\times C.$
Clearly, $p_{1}\circ p_{12}=\text{p}_{1}$.
Assume we have fixed an action of $\mathscr{O}(X)$ on $\mathscr{O}(C)$ (i.e., a coaction scheme morphism $C\to X\times C$). Suppose that $\Phi\in C^{3}(X,\mathbb{G}_{m})$ and $\gamma\in C^{2}(C,\mathbb{G}_{m})$ are cochains 222 I.e., $\gamma\in\mathscr{O}(C)^{\otimes 2}$, $\Phi\in\mathscr{O}(X)^{\otimes 3}$ are invertible, $(\varepsilon\otimes{\rm id})(\gamma)=({\rm id}\otimes\varepsilon)(\gamma)=1$ and $(\varepsilon\otimes{\rm id}\otimes{\rm id})(\Phi)=({\rm id}\otimes\varepsilon\otimes{\rm id})(\Phi)=({\rm id}\otimes{\rm id}\otimes\varepsilon)(\Phi)=1$. such that
@@
    ${p_{1}^{*}S\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{({\rm id}\times{\rm m})^{*}(\rho)}$$\scriptstyle{p_{12}^{*}(\rho)}$${(\mu\circ p_{12})^{*}S\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(\mu\times{\rm id})^{*}(\rho)}$${\nu^{*}S\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi\cdot({\rm id}\otimes{\rm id}\otimes\gamma)}$${\nu^{*}S}$
@@
Let $\gamma\in Z^{2}(C,\mathbb{G}_{m})$, $C_{\gamma}$ the finite group scheme central extension of $C$ by $\mathbb{G}_{m}$ associated with $\gamma$, and ${\rm Rep}_{k}(C,\gamma)$ the category of finite dimensional representations of $C_{\gamma}$ on which $\mathbb{G}_{m}$ acts trivially. Then ${\rm Rep}_{k}(C,\gamma)\cong{\rm Coh}^{(C,\gamma)}({\rm pt})$ as abelian categories. ∎
@@
[…]
@@
Retain the setting of §2.4. Let $\mathscr{O}(C)_{\gamma}$ denote the vector space $\mathscr{O}(C)$ equipped with the $k$-linear map $\Delta_{\gamma}$ given by $\Delta_{\gamma}(f):=\Delta(f)\gamma$, where $\Delta$ is the standard comultiplication map of $\mathscr{O}(C)$.
@@
$(\Delta_{\gamma}\otimes{\rm id})\Delta_{\gamma}=\Phi\cdot({\rm id}\otimes\Delta_{\gamma})\Delta_{\gamma}.$
@@
    ${\rm Coh}^{(C,\gamma)}(X,\Phi)\cong{\rm Comod}(\mathscr{O}(C)_{\gamma})_{{\rm Coh}(X,\Phi)},$
    ${}^{(B,\beta)}{\rm Coh}(X,\Phi)\cong{\rm Comod}_{{\rm Coh}(X,\Phi)}(\mathscr{O}(B)_{\beta}),\,\,\,{\rm and}\,\,\,$
    ${\rm Coh}^{((B,\beta),(C,\gamma))}(X,\Phi)\cong{\rm Bicomod}_{{\rm Coh}(X,\Phi)}(\mathscr{O}(B)_{\beta},\mathscr{O}(C)_{\gamma}).$
    In particular, the categories ${\rm Coh}^{(C,\gamma)}(X,\Phi)$, ${}^{(B,\beta)}{\rm Coh}(X,\Phi)$ and ${\rm Coh}^{((B,\beta),(C,\gamma))}(X,\Phi)$ are abelian.
@@
    If $\mathscr{I}\subset\mathscr{O}(X)$ is a $C$-stable ideal ($B$-stable, $(B,C)$-bistable, respectively), then for any $S\in{\rm Coh}^{(C,\gamma)}(X,\Phi)$, $\mathscr{I}S$ is a subobject of $S$ in ${\rm Coh}^{(C,\gamma)}(X,\Phi)$ ${}^{(B,\beta)}{\rm Coh}(X,\Phi)$, ${\rm Coh}^{((B,\beta),(C,\gamma))}(X,\Phi)$, respectively).
@@
2.6. (Bi)equivariant morphisms
Retain the setting of §2.4. Assume that $Y$ is a finite scheme on which $B$ acts from the left via $\mu_{B\times Y}$ and $C$ acts from the right via $\mu_{Y\times C}$.
Proposition 2.13.
The following hold:
  1. (1)
  2. (2)
  3. (3)
Proof.
We prove (1), the proofs of (2) and (3) being similar.
Finally, it is straightforward to verify that the adjunction
as desired. ∎
Example 2.14.
2.7. The tensor category ${\rm Coh}(G,\omega)$
${\rm Coh}(G,\omega)\cong{\rm Rep}_{k}(\mathscr{O}(G),\omega)$
For any closed point $g\in G(k)$, let
$\omega_{3}^{g}:=({\rm Ad}g\otimes{\rm Ad}g)({\rm id}\otimes{\rm id}\otimes g)(\omega),$
$\omega_{1}:=(g\otimes{\rm id}\otimes{\rm id})(\omega),\,\,\,{\rm and}\,\,\,\omega_{2}:=({\rm id}\otimes g\otimes{\rm id})(\omega),$
and set
For example, in the étale case, we have

### v2

@@ Abstract.
Let ${\mathscr{C}}:={\mathscr{C}}(G,\omega,H,\psi)$ be a finite group scheme-theoretical category over an algebraically closed field of characteristic $p\geq 0$ [G1]. For any indecomposable exact module category over $\mathscr{C}$}, we classify its simple objects and provide an expression for their projective covers in terms of double cosets and projective representations of certain closed subgroup schemes of $G$. This upgrades a result of Ostrik [O] for group-theoretical fusion categories in characteristic 0, and generalizes our previous work [GS] for the case $\omega=1$. As a byproduct, we describe the simples and indecomposable projectives of $\mathscr{C}$}. Finally, we apply our results to describe the blocks of the center of ${\rm Coh}(G,\omega)$.
@@ 1. Introduction
Recall that group-theoretical fusion categories [O] play a fundamental role in the theory of fusion categories in characteristic 0 (see [GS, Section 1] and references therein), which suggests that finite group scheme-theoretical categories in characteristic $p>0$ play an important role in the theory of finite tensor categories, thus making them interesting tensor categories worthwhile to study.
@@
Unlike in characteristic 0, finite group scheme-theoretical categories and their module categories are rarely semisimple if $p>0$, and a plethora of questions arise. In [GS], we addressed some of these questions in the case $\omega=1$, and the goal of this work is to extend that treatment to arbitrary $\omega\in Z^{3}(G,\mathbb{G}_{m})$.
@@
${\rm Fun}_{{\rm Coh}(G,\omega)}\left(\mathscr{M}(H,\psi),\mathscr{M}(K,\eta)\right)\simeq{\mathscr{M}}\left((H,\psi),(K,\eta)\right),$
@@
Finally, define the abelian categories
$\mathscr{A}:=\mathrm{Coh}^{\left(\left(B,\psi\right),\left(C,\eta\right)\right)}(A\backslash(B\times C),\Phi),$
$\mathscr{B}:=\mathrm{Coh}^{\left(\left(A\times B,\xi^{-1}\times\psi\right),\left(C,\eta\right)\right)}(B\times C,\Phi)$
(see Definition 2.7), and let
${\rm Rep}(A,\xi^{-1})_{k}:={\rm Corep}_{k}(\mathscr{O}(A)_{\xi^{-1}}),$
${\rm Rep}_{k}(A,\xi):={\rm Corep}(\mathscr{O}(A)_{\xi})_{k}$
be the categories of finite dimensional left, right comodules over the twisted coalgebras $\mathscr{O}(A)_{\xi^{-1}}$, $\mathscr{O}(A)_{\xi}$, respectively. By (2.33), there is a canonical equivalence of categories
${\rm Rep}(A,\xi^{-1})_{k}\simeq{\rm Rep}_{k}(A,\xi).$
@@
${\rm Rep}(A,\xi^{-1})_{k}$$\mathscr{B}$$\mathscr{A}$$\scriptstyle{{\rm Ind}_{(A,\xi^{-1})}^{\left(\left(B,\psi\right),\left(C,\eta\right)\right)}}$$\scriptstyle{{\rm Res}_{(A,\xi^{-1})}^{\left(\left(B,\psi\right),\left(C,\eta\right)\right)}}$
@@
${\rm F}\colon{\rm Rep}(A,\xi^{-1})_{k}\xrightarrow{\simeq}\mathrm{Coh}^{\left(\left(B,\psi\right),\left(C,\eta\right)\right)}(A\backslash(B\times C),\Phi),$
@@
    $\mathbf{Ind}_{Z}:{\rm Rep}(L^{g},\xi_{g}^{-1})_{k}\xrightarrow{\simeq}{\mathscr{M}}_{Z},$
@@
    There is a bijection between equivalence classes of pairs $(Z,V)$, where $Z\in Y(k)$ is a closed point with representative $g\in Z(k)$, and $V\in{\rm Rep}(L^{g},\xi_{g}^{-1})_{k}$ is simple, and simple objects of $\mathscr{M}$, assigning $(Z,V)$ to $\mathbf{Ind}_{Z}(V)$.
@@
    $\mathbf{F}_{Z}:{\rm Rep}(L^{g},\xi_{g}^{-1})_{k}\xrightarrow{\simeq}{\mathscr{M}}_{Z}.$
@@
$\mathbf{F}_{C}:{\rm Rep}_{k}(G_{C},\omega_{g})\xrightarrow{\simeq}\mathscr{Z}(G,\omega)_{C},\,\,\,V\mapsto\mathscr{O}(C)\otimes_{k}V,$
@@
$F:\mathscr{Z}(G,\omega)\xrightarrow{\simeq}\mathscr{Z}_{\mathscr{D}^{\circ}}\left(\mathscr{D}\right)^{G(k)}=\left(\bigoplus_{a\in G(k)}\mathscr{Z}_{\mathscr{D}^{\circ}}\left(\mathscr{D}^{\circ}\boxtimes a\right)\right)^{G(k)}.$
@@
$F_{C}:\overline{\mathscr{Z}(G,\omega)_{C}}\xrightarrow{\simeq}\bigoplus_{a\in C(k)}\mathscr{Z}_{\mathscr{D}^{\circ}}\left(\mathscr{D}^{\circ}\boxtimes a\right)^{G(k)}.$
@@
$F_{1}:\overline{{\rm Rep}_{k}(G)}\xrightarrow{\simeq}\mathscr{Z}\left(G^{\circ},\omega^{\circ}\right)^{G(k)}.$
@@ 2. Preliminaries
2.1. Conventions
We work over an algebraically closed field $k$ of characteristic $p\geq 0$.
All schemes $X$ considered in this paper are assumed to be finite over $k$, and equipped with a scheme morphism ${\rm Spec}(k)\to X$, unless otherwise stated. Equivalently, we will assume that $\mathscr{O}(X)$ is a finite dimensional commutative $k$-algebra with augmentation map $\varepsilon:\mathscr{O}(X)\to k$.
We assume familiarity with the theory of finite tensor categories and their modules categories, and refer to [EGNO] for any unexplained notion.
@@ 2.2. Sheaves on finite schemes
For a finite scheme $X$ as in §2.1, let ${\rm Coh}(X)$ be the abelian category of sheaves on $X$, i.e., the category of finite dimensional representations of the finite dimensional commutative algebra $\mathscr{O}(X)$. Given a closed point $x\in X(k)$, let ${\rm Coh}(X)_{x}\subset{\rm Coh}(X)$ be the abelian subcategory of sheaves on $X$ supported on $x$, so
@@
Let $Y$ be a finite scheme as in §2.1, $\varphi:Y\to X$ a scheme morphism, and $\varphi^{\sharp}:\mathscr{O}(X)\to\mathscr{O}(Y)$ the corresponding algebra homomorphism. Recall that $\varphi$ induces a pair $(\varphi^{*},\varphi_{*})$ of adjoint functors of abelian categories
@@ 2.3. Finite group schemes
A finite group scheme $G$ is a finite scheme whose coordinate algebra $\mathscr{O}(G)$ is a (finite dimensional commutative) Hopf algebra (see, e.g., [J, W]). It is called étale if $G=G(k)$, where $G(k)$ is the group of closed points of $G$; these finite group schemes are in correspondence with finite abstract groups (see, e.g., [W, Section 6.4]). On the other hand, $G$ is connected if $G(k)=1$. For example, if $\mathfrak{g}$ is a finite dimensional $p$-Lie algebra over $k$ ($p>0$), its $p$-restricted universal enveloping algebra $u(\mathfrak{g})$ is a finite dimensional cocommutative Hopf algebra (see, e.g., [J]), and the dual commutative Hopf algebra $u(\mathfrak{g})^{*}$ is local, and satisfies $x^{p}=0$ for all $x$ in the augmentation ideal. Thus, the associated group scheme is connected. Moreover, any finite connected group scheme is obtained by successive extensions of dualized $p$-restricted enveloping algebras (see [DG, Section 2.7]).
@@
(2.3) $1\to G^{\circ}\xrightarrow{i}G\mathrel{\mathop{\rightleftarrows}^{\pi}_{q}}G(k)\to 1,$
where $G^{\circ}$ is connected and $G(k)$ is étale. In particular, $G(k)$ acts on ${\rm Coh}(G)_{1}\cong{\rm Coh}(G^{\circ})$, say via $a\mapsto T_{a}$, and we have an equivalence
(2.4) ${\rm Coh}(G)\simeq{\rm Coh}(G^{\circ})\rtimes G(k).$
@@ 2.4. Quotients and free actions
Let $C$ be a finite group scheme, and $X$ a finite scheme as in §2.1, equipped with a right $C$-action
@@
is a closed immersion, where $p_{1}:X\times C\to X$ is the obvious projection morphism. Recall [Mu, Theorem 1(B), p.112] that in this case, the morphism $(p_{1},\mu)$ induces a scheme isomorphism
@@
[…]
@@
Consider the obvious projection morphisms
$p_{1}:X\times C\twoheadrightarrow X,\,\,\,q_{1}:X\times C\times C\twoheadrightarrow X,$
${\rm and}\,\,\,q_{12}:X\times C\times C\twoheadrightarrow X\times C.$
Clearly, $p_{1}\circ q_{12}=q_{1}$.
Assume we have fixed an action of $\mathscr{O}(X)$ on $\mathscr{O}(C)$ (i.e., a coaction scheme morphism $C\to X\times C$). Suppose that $\Phi\in C^{3}(X,\mathbb{G}_{m})$ and $\gamma\in C^{2}(C,\mathbb{G}_{m})$ are cochains 222 I.e., $\gamma\in\mathscr{O}(C)^{\otimes 2}$, $\Phi\in\mathscr{O}(X)^{\otimes 3}$ are invertible, $(\varepsilon\otimes{\rm id})(\gamma)=({\rm id}\otimes\varepsilon)(\gamma)=1$ and $(\varepsilon\otimes{\rm id}\otimes{\rm id})(\Phi)=({\rm id}\otimes\varepsilon\otimes{\rm id})(\Phi)=({\rm id}\otimes{\rm id}\otimes\varepsilon)(\Phi)=1\otimes 1$. such that
@@
    ${q_{1}^{*}S\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{({\rm id}\times{\rm m})^{*}(\rho)}$$\scriptstyle{q_{12}^{*}(\rho)}$${(\mu\circ q_{12})^{*}S\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(\mu\times{\rm id})^{*}(\rho)}$${\nu^{*}S\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi\cdot({\rm id}\otimes\gamma)}$${\nu^{*}S}$
@@
Let $\gamma\in Z^{2}(C,\mathbb{G}_{m})$, $C_{\gamma}$ the finite group scheme central extension of $C$ by $\mathbb{G}_{m}$ associated with $\gamma$, and ${\rm Rep}_{k}(C,\gamma)$ the category of finite dimensional left representations of $C_{\gamma}$ on which $\mathbb{G}_{m}$ acts trivially. Then ${\rm Rep}_{k}(C,\gamma)\cong{\rm Coh}^{(C,\gamma)}({\rm pt})$ as abelian categories. ∎
@@
${B\times X\times C\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\rm id}\otimes\mu_{X\times C}}$$\scriptstyle{(\mu_{B\times X}\otimes{\rm id})\Phi}$${X\times C\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu_{X\times C}}$${B\times X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu_{B\times X}}$${X.}$
@@
2.6. (Bi)equivariant morphisms
Retain the setting of §2.5. Assume that $Y$ is a finite scheme as in §2.1, on which $B$ acts from the left via $\mu_{B\times Y}$ and $C$ acts from the right via $\mu_{Y\times C}$.
Proposition 2.8.
The following hold:
  1. (1)
  2. (2)
  3. (3)
Proof.
We prove (1); the proofs of (2) and (3) being similar.
Finally, it is straightforward to verify that the adjunction
as desired. ∎
Example 2.9.
2.7. The tensor category ${\rm Coh}(G,\omega)$
${\rm Coh}(G,\omega)\simeq{\rm Rep}_{k}(\mathscr{O}(G),\omega)$
For any closed point $g\in G(k)$, let
(2.13) $\omega_{3}^{g}:=({\rm Ad}g\otimes{\rm Ad}g)({\rm id}\otimes{\rm id}\otimes g)(\omega),$
(2.14) $\omega_{1}:=(g\otimes{\rm id}\otimes{\rm id})(\omega),\,\,\,{\rm and}\,\,\,\omega_{2}:=({\rm id}\otimes g\otimes{\rm id})(\omega),$
and set
For example, in the étale case, we have
@@ 2.7. The tensor category ${\rm Coh}(G,\omega)$
Retain the setting of §2.5. Assume further that $X$ is a finite group scheme, and $\Phi\in\mathscr{O}(X)^{\otimes 3}$ is a Drinfeld associator, so that ${\rm Coh}(X,\Phi)$ is a finite tensor category (see §2.7).
Let $\mathscr{O}(C)_{\gamma}$ denote the vector space $\mathscr{O}(C)$ equipped with the $k$-linear map $\Delta_{\gamma}$ given by $\Delta_{\gamma}(f):=\Delta(f)\gamma$, where $\Delta$ is the standard comultiplication map of $\mathscr{O}(C)$.
@@
$\Phi\cdot(\Delta_{\gamma}\otimes{\rm id})\Delta_{\gamma}=({\rm id}\otimes\Delta_{\gamma})\Delta_{\gamma}.$
@@
    ${\rm Coh}^{(C,\gamma)}(X,\Phi)\simeq{\rm Comod}(\mathscr{O}(C)_{\gamma})_{{\rm Coh}(X,\Phi)},$
    ${}^{(B,\beta)}{\rm Coh}(X,\Phi)\simeq{\rm Comod}_{{\rm Coh}(X,\Phi)}(\mathscr{O}(B)_{\beta}),\,\,\,{\rm and}\,\,\,$
    ${\rm Coh}^{((B,\beta),(C,\gamma))}(X,\Phi)\simeq{\rm Bicomod}_{{\rm Coh}(X,\Phi)}(\mathscr{O}(B)_{\beta},\mathscr{O}(C)_{\gamma}).$
    In particular, the categories ${\rm Coh}^{(C,\gamma)}(X,\Phi)$, $^{(B,\beta)}{\rm Coh}(X,\Phi)$, and ${\rm Coh}^{((B,\beta),(C,\gamma))}(X,\Phi)$ are abelian.
@@
    If $\mathscr{I}\subset\mathscr{O}(X)$ is a $C$-stable ideal ($B$-stable, $(B,C)$-bistable, respectively), then for any $\in{\rm Coh}^{(C,\gamma)}(X,\Phi), S\mathscr{I}S$ is a subobject of $S$ in ${\rm Coh}^{(C,\gamma)}(X,\Phi)$ ($^{(B,\beta)}{\rm Coh}(X,\Phi)$, ${\rm Coh}^{((B,\beta),(C,\gamma))}(X,\Phi)$, respectively).